OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side of the right triangle with the given measures. c is the hypotenuse. a = 6, b = 10, c = ? Answer choices: 11.7 8 136

OpenStudy (anonymous):

Pythagorean therorem a^2 + b^2 = c^2 6^2 + 10^2 = c^2 or c = sqrt( 6^2 + 10^2)

OpenStudy (kymber):

Plug in & solve